OpenStudy (anonymous):

solve the system by the substitution method 3x-9y=3 -2x+6y=5

OpenStudy (pinklion23):

are you looking for x

OpenStudy (pinklion23):

so for the first one divide both sides by 3

OpenStudy (pinklion23):

then add 9 to both sides

OpenStudy (pinklion23):

so for the first one x=9

OpenStudy (pinklion23):

for the second one divide both sides by -2

OpenStudy (pinklion23):

then subtract 6

OpenStudy (pinklion23):

then you have xy= -8.5

OpenStudy (anonymous):

System has no solution; because ich you multiply first by 2 and second equation by 3 you get: 6x-18y = 6 -6x +18y = 15 after add together you get: 0= 6+15 So It's not possible -> equations has no solutions Also checkout machine solution: http://www.equationcalculator.org/?input=3x-9y%3D3%0D%0A-2x%2B6y%3D5&submit=Calculate